OpenStudy (anonymous):

Which sentence explains the correct first step in the solution of this equation? 4 • (x – 3) = 9 Add 2 to both sides. Apply the distributive property to get 4x – 12 = 9. Apply the distributive property to get 4x – 3 = 9. Subtract 9 from both sides.

OpenStudy (anonymous):

Is it 4*(x-3)=9 or 4-(x-3)=9 ?

OpenStudy (anonymous):

4 • (x – 3) = 9

OpenStudy (anonymous):

@Andras

OpenStudy (anonymous):

First thing to do 99% of the time is the get rid of the brackets. Or using proper maths applying the distributive property

OpenStudy (anonymous):

So 4*(x-3)=??

OpenStudy (anonymous):

Can you expand the bracket? Whatn you need to do is multiply by 4 both terms inside the bracket